OpenStudy (anonymous):

simplify the radical expression square root of 63x^15y^9/7xy^11

OpenStudy (anonymous1):

\[\sqrt{ \frac{63x^{15}y^9} {7xy^{11}} }\] \[=\sqrt{ \frac{63x^{14}} {7y^{2}} }= \sqrt{\frac{63}{7}}\frac{x^7}{y}=\sqrt{9}\frac{x^7}{y}\] \[=3\frac{x^7}{y}\]

OpenStudy (anonymous1):

So, the final answer is \[3\frac{x^7}{y}\]